. On the day Saffron was born, her parents put $800 into an investment account that promises to pay a fixed interest rate of 7 percent per year. How much money will she have in this account when she turns 18? Round to two decimal places. 2. At what rate must $311.25 be compounded annually for it to grow to $612.28 in 10 years? Submit your answer as a percentage and round to two decimal places. How much money must be put into a bank account yielding 2.50% (compounded annually) in order to have $1,875 at the end of 13 years? Round to two decimal places 3, Yolanda deposited $8,000 in a bank account, and 9 years later she closes out the account, which is worth $16,000. What annual rate of interest has she earned over the 9 years? Submit your answer as a percentage and round to two decimal places 4. Bridgette wants to retire 36 years from now. She decides to start saving $400 each month into a Roth IRA starting at the end of this month. If the IRA is expected to earn an average return of 9% annually, how much will she have in the account at the end of 5. 36 years? Round to two decimal places 6. Durran has recently acquired a rare art piece that he plans to put on display in his private collection. He estimates that revenues generated from donations and admissions tickets to see the new exhibit will be $2,000 per month for the next eight years. If he requires a rate of return of 12%, how much are the expected cash flows worth for him today? Round to two decimal places. Adelai Inc. loans you $32,000 for five years to buy a car. The loan must be repaid in 60 equal monthly payments. The annual interest rate on the loan is 8percent. What is your monthly payment? Round to two decimal place 7. 8. Your company has received a $50.000 loan from an industrial finance company. The annual payments are $6,202.70. If the company is paying 9 percent interest per year, how many l oan payments must the company make? Round to the nearest number of periods
Answer #1

As per rules I am answering the first 4 subparts of the question

1: FV = PV*(1+r)^n

PV = 800

R= 7%

N = 18

Fv = 800*(1+0.07)^18 = $2703.95

2: FV = PV*(1+r)^n

612.28 = 311.25*(1+r)^10

(612.28/ 311.25)^(1/10) = 1+r

R= 7%

3: PV = FV/(1+r)^n

=1875/ (1+0.025)^13

=$1360.16

4: FV = PV*(1+r)^n

16000= 8000*(1+r)9

(16000/8000)^(1/9) = 1+r

r= 8.006%